# Is Human-Like Text Liked by Humans? Multilingual Human Detection and Preference Against AI

### v1

@@ 1 Introduction
Recent technological developments have advanced the generative artificial intelligence (AI) models, such as GPT-*, Gemini, Claude, and Llama OpenAI (2023); Team et al. (2023); Anthropic (2024); Dubey et al. (2024), thus aggressively blurring the lines between human and AI capabilities. How often can state-of-the-art (SOTA) LLMs fool human evaluators, i.e., passing the Turing test? Can human evaluators correctly predict the origins of encountered content (human vs. machine), thus safeguarding against the potential misuse of LLMs? Several studies have explored whether humans can distinguish between content generated by a human vs. a machine. Unsurprisingly, the findings depend in part on the quality of the machine generator. Early studies probing the outputs of less advanced AI models found that human evaluators could indeed detect the difference van der Lee et al. (2019). However, in studies using LLMs such as GPT-3.5-turbo, human evaluators are frequently close to random chance Guo et al. (2023); Hitsuwari et al. (2023); Dugan et al. (2023); Chein et al. (2024); Wang et al. (2024a), particularly for laymen who are not frequent users of LLMs. While covering different domains, e.g., academic paper abstracts, Wikipedia paragraphs, question-answering responses Wang et al. (2024b, a), news and review comments Chein et al. (2024), most studies focused on GPT-3.5-turbo and English (see Table 1). Generally, fewer than 300 examples were evaluated by experts (LLM researchers or frequent users) in such studies. With the advancement of newer LLMs such as GPT-4o, Claude-3.5-Sonnet, and Llama3.1, we ask whether these findings generalize to them, to other languages, and to native expert annotators? What are the upper bounds of human detection performance across languages, domains and new LLMs? To answer these questions, we perform a comprehensive case study over sixteen datasets spanning nine languages, nine domains, and eleven SOTA LLMs. We specifically focus on the following four research questions: (iHow well can human annotators distinguish human-written text from SOTA LLMs generations? (iiWhat are the notably detectable linguistic qualities in human- and AI-authored texts that shape decisions about text origin? (iiiCan a prompting strategy fill the gap between human and machine? and (ivDo humans always prefer human-authored text? Our human preference ranking dataset can serve as a useful alignment source for multilingual LLM improvement. Our comprehensive analysis over both detection and analysis exposes the major gaps between current LLMs and human expectations, thus suggesting possible directions for future LLM development. Moreover, our human detection evaluation challenges previous findings in terms of random-guess detection capability for humans. Our contributions are as follows:
@@
  • Our extensive case study shows that expert annotators achieve detection accuracy of 87.6% on average over $\sim$9K examples, demonstrating that humans can identify LLM outputs and safeguard against LLM misuse. We identified a spectrum of language- and domain-specific distinguishable features between a human and a machine, and we summarize five generalized distinctions across languages, which clearly expose the limitations of current LLM outputs and indicate directions for improvements.
@@
  • Humans do not always like human-written text: preferences are evenly split between human vs. machine, with a strong tendency to favor machine-generated text (MGT) in Russian and Arabic (for domains such as summary and tweets).
  • We release 17K original and 32K improved MGTs, labels of 13.5K human detection instances, 5K preferences and 1.6K fill-in-the-gap survey, and metadata for 19 annotators to facilitate future work on human vs. MGT detection, and their relationship with individual characteristics and preferences.
@@
Guo et al. (2023) 2-class English GPT-3.5 Wiki text, QA 150 8 experts + 9 layman 48-90%
Guo et al. (2023) 2-class Chinese GPT-3.5 Wiki text, QA 210 8 experts + 9 layman 54-93%
@@
Таблица 1: Experts refers to individuals who are frequent users of LLMs, laymen are people who have never heard of or used LLMs or have used them only rarely, AMT is Amazon Mechanical Turk.
@@ 2 Case Study
Previous studies presented the difficulty humans face in distinguishing SOTA LLM-generated content from human-written text, often resulting in a random guess (see more in Appendix A). However, most evaluations focused on English and generations by GPT-3.5-turbo, leaving the detectability of MGT in other languages and LLMs uncertain. To verify whether this observation can be generalized to other languages and more advanced LLMs, we collected LLM generations based on 16 datasets across nine domains and nine languages. 19 native speakers who are either LLM researchers or practitioners performed human evaluations, investigating (i) whether humans can correctly discern human vs. AI outputs, and (ii) whether humans prefer fellow human answers or LLM responses.
@@ 2.1 Task and Dataset
The goal is to identify whether the text was written by a human or generated by models given a single text, or to recognize which text is written by a human given a pair of texts: one human-written and one machine-generated. In data collection, we focused on datasets from common domains such as community QA, news, tweets, and government reports, alongside domains requiring high-integrity LLM applications, including educational and academic contexts, such as accurate knowledge verification in Wikipedia-style texts, and identifying the authorship of student essays and peer reviews. For a given language and given a dataset, we sampled 300-600 human texts and then generated corresponding machine text using two SOTA LLMs: a multilingual model (e.g., from the GPT or the Claude series) and a language-specific model (e.g., ChatGLM or Qwen for Chinese and AceGPT for Arabic), to analyze the impact of different LLMs on detection performance, particularly for non-English languages. As shown in Table 4, we collected data based on 16 datasets across nine languages. The generation prompts and collection details are shown in Appendix B.
2.2 Human Detection Setups
Annotation Settings
To mimic real-world machine-generated text detection scenarios, we set up four human evaluation settings. Given human-written text and machine-generated text, representing by $hwt$ and $mgt$ respectively, human annotators are asked to identify which text was written by a human. Note that $mgt$ can be generated by multiple different LLMs, referring to as $mgt_{i}$, where $i\in[1,2,\cdots,n]$. As shown in Table 2, according to the input and the output options, we categorize detection settings as I. pair-binary, II. pair-four-class, III. single-binary, and IV. triplet-three-class. For single text input, either $hwt$ or $mgt$, the goal is to recognize whether the text was written by a human, by answering just Yes or No. This is suitable for the scenario where for the human text there is no necessarily a corresponding machine-generated text, and thus they can be collected from different sources and for different topics, such as Arabic tweets. Given a pair of texts (text1, text2), a binary output is easier than the four-class detection. The pair-binary setting asks that either text1 or text2 is $hwt$, and the other one is $mgt$, while the pair-four-class setting has no restrictions: each of text1 and text2 can be $hwt$ or $mgt$, regardless of the label of the other text. Sometimes, we want to compare human text to generations from different LLMs, in which case, we apply IV, which we limit to a three-class detection: human vs. LLM1 vs. LLM2. Overall, I and IV are suitable for scenarios where there is a human text and its corresponding machine-generated text. II and III can be used if there are non-corresponding human-written and machine-generated texts. If the annotators have seen some human-written and some machine-generated text before detection, we refer to this as a few-shot setting; otherwise, we have zero-shot.
@@ 2.2 Human Detection Setups
IV. Triplet-Three-Class ($hwt$, $mgt_{1}$, $mgt_{2}$) Given a set of texts (text1, text2, text3), identify which one is human-written? One of the text1, text2 and text3 must be $hwt$, and others are $mgt$ randomly sampled from $mgt_{i}$. A. text1; B. text2; C. text3 parallel human and multiple LLM generations are collected to make comparisons.
@@
Annotation Tool
To mitigate potential labeling biases arising from raw spreadsheet annotation and to enhance efficiency, we implemented two methods with optimized interfaces and workflows for our annotation: (1) a custom pipeline using the Google Workspace suite, including Apps Script, Google Sheets, and Google Forms. The core idea was to store all data in Google Sheets, use Apps Script to extract data and generate a survey in Google Forms; and (2) Label Studio, an open-source multi-type data labeling and annotation tool with a standardized output format. We designed a custom template for our annotation task and collected results using this platform. The annotators were given the choice to use their preferred tool.
@@
In order to explore the upper bound of human detection capability, instead of using crowd-sourcing annotators, we conducted in-house labeling. The annotators were BSc, MSc, and PhD students, as well as postdocs, who were familiar with NLP tasks and LLM generations. All annotators independently completed their individual annotations. For each language, the annotators were all native speakers of that language. See more detail in Appendix A.3.
3 Human Detection
We performed an extensive case study on 9K instances across nine languages to verify how difficult it is for native speakers to detect AI outputs in everyday domains. Table 3 demonstrates that the average human detection accuracy is 87.6%. This reveals that this is not particularly difficult for native human experts, contrary to what previous studies have reported. Below, we zoom into the impact of various factors.
@@ 3 Human Detection
Таблица 3: Human annotator detection accuracy over 16 datasets and 9 languages: we have a total of 30 annotation settings and 19 unique human annotators. The average accuracy of the human expert guesses is 87.6%.
@@ 3.3 Generator
It is hard to detect MGT across generators and languages. While there are minimal differences for Chinese (accuracy is 100% vs. 98% for GPT-4o vs. Qwen-turbo), there are sizable differences for Italian and Arabic. Based on Italian DICE News, the same annotator detected generations by Anita (an Italian fine-tuned Llama3-8B), Llama3-405B, and GPT-4o, achieving accuracy of 88%, 99%, and 100%, respectively. Similarly, for Arabic tweets, GPT-4o’s outputs are more similar to human text and thus more difficult to detect compared to those by Qwen2, as shown in Table 6.
@@ 3.4 Annotation Setting
We conducted the majority of our annotations under setting I. pair-binary: given a pair ($hwt$, $mgt$), it asks to identify which of the two texts is human-written. We assumed that more complex settings would be more challenging. For instance, II. pair-four-class should be harder, as each of the texts could be human- or machine-generated, independently of the other. Yet, for Chinese student essays, the performance for II does not degrade compared to I. Similarly, for government reports in IV. triplet-three-class, where the annotators have to select the human-written text among three candidates, there was no degradation compared to I. However, III. single-binary proves to be more difficult than I. pair-binary for both Arabic and Russian. While domain differences do exist, e.g., tweets vs. summary vs. news in Arabic and news vs. summary for Russian, the substantial performance gap (>20%) can still be partially attributed to the annotation settings. Overall, comparing the results for I. vs. III., it is easier to distinguish machine-generated content when given a comparison pair, rather than for single answer. Yet, introducing text triplets or increasing the number of machine-generated or human-written texts had minimal impact on detection performance. Moreover, using few shots before detection boosted the confidence of the annotators, resulting in higher accuracy compared to zero-shot. For datasets with high accuracy, before seeing labeled samples, the annotators found the distinction to be obvious. After seeing a few examples, the annotator was extremely confident in distinguishing human vs. machine text based on indicative features of MGT.
@@ 3.6 Distinguishable Factors
We summarize five remarkable distinguishable signals between machine-generated vs. human-written text across the 16 datasets and the 9 languages; see more details in Appendix C.
@@
  • Machine-generated text shows a mixture of other languages. Non-English language responses often contain some English parts, which is very rare in human text.
@@ 4 Can Prompting Fill in the Gap?
Given that LLMs can strictly follow instructions and their outputs are heavily influenced by the system and the user prompts, we investigated whether explicitly instructing LLMs to mimic human style can help narrow the gap. Responding to the distinguishable factors summarized in Appendix C for each dataset, we asked the human annotators to craft new prompts, aiming to improve the generations and to reduce the gap between human-written and LLM-generated texts. This involved trying instructions that (1) incorporate specific details and references, (2) avoid formulaic structures and formats, e.g., bullet points and Markdown, and (3) generate texts of varying length, structure, and sentiment. Table 8 presents the results for both the original and the improved prompts for all datasets. Measurements: We re-generated the machine-generated parts of the text pairs, using the same models with improved prompts, and then sampled 200–600 examples from each dataset to assess whether and to what extent, the prompting strategy narrowed the gap between human-written and machine-generated texts. We used two approaches: (1) fill-the-gap survey, where the original annotators evaluated whether the newly-generated text bridged the gap (Yes, No, or Partially), and (2) downstream detection, where we compared the detection accuracy before and after applying the improved prompts. A decline in detection accuracy indicated a reduced distinction between human and machine text, making the differentiation more difficult, and further revealing that prompting was effective. Our experiments involved both human annotator evaluation and automated detection.
@@
The original annotators who conducted the detection on earlier generations were asked to evaluate whether the new prompts addressed the gaps for each example: Yes, No, and Partially. The distributions across the six datasets in Figure 1 shows that in about 50% of the cases, the prompt adjustments were effective to either fully or partially mitigate the gaps. Large improvements were observed for Kazakh Wikipedia and Arabic tweets. For the former, the revised prompt reduced repetitive sentence patterns (more diverse), but the formulaic expressions were not entirely eliminated. New outputs also included more concrete information, such as dates and names, while the inclusion of culturally-nuanced details remained challenging. The newly-generated Arabic tweets could touch on relatable human topics and express genuine emotions tied to daily experiences; however, the frequently added irrelevant hashtags at the end of the tweets made them easily identifiable as machine-generated. Moreover, the tweets often leaned on an overly optimistic tone even when negative experiences were mentioned. The annotators for English peer reviews noted that, despite the prompt adjustments, the models remained highly formulaic in their outputs, the length of the reviews remained relatively uniform, and the overall structure lacked variance. This may be due to the inherent nature of the peer review domain, while human reviews exhibited more variability in both length and structure, feeling more organic. Similar issues remained for Chinese student essays (formulaic structure by using ‘‘firstly, then, moreover, finally’’ persisted) and government reports (certain repetitive phrases). Overall, adjusting the prompts did fill some gaps, but cultural nuances, diversity of length, structure and phrases, and sentiment adaption to scenarios remained challenging. See more in Appendix D. Since annotators’ background may influence the survey results (see D.10), we conducted a second round of MGT detection on the new generations with the hop e for more objective results.
Рис. 2: Human detection accuracy for the original vs. the improved generations.

### v2

@@ 1 Introduction
Recent technological developments have advanced the generative artificial intelligence (AI) models, such as GPT-*, Gemini, Claude, and Llama OpenAI (2023); Team et al. (2023); Anthropic (2024); Dubey et al. (2024), thus aggressively blurring the lines between human and AI capabilities. How often can state-of-the-art (SOTA) LLMs fool human evaluators, i.e., passing the Turing test? Can human evaluators correctly predict the origins of encountered content (human vs. machine), safeguarding against the potential misuse of LLMs? Several studies have explored whether humans can distinguish between content generated by a human vs. a machine. Unsurprisingly, the findings depend in part on the capability of generative models. Early studies probing the outputs of less advanced AI models found that human evaluators could indeed detect the difference van der Lee et al. (2019). However, in studies using LLMs like GPT-3.5-turbo, human evaluators are frequently close to random chance Guo et al. (2023); Hitsuwari et al. (2023); Dugan et al. (2023); Chein et al. (2024); Wang et al. (2024a), particularly for laymen who are not frequent users of LLMs. While covering different domains, e.g., academic paper abstracts, Wikipedia paragraphs, question-answering responses Wang et al. (2024b, a), news and review comments Chein et al. (2024), most studies focused on GPT-3.5-turbo and English texts (see Table 1). Generally, fewer than 300 examples were evaluated by experts (LLM researchers or frequent users) in such studies. With the advancement of newer LLMs such as GPT-4o, Claude-3.5-Sonnet, and Llama4, we ask whether these findings generalize to them, to other languages, and to native expert annotators? What are the upper bounds of human detection performance across languages, domains and new LLMs? To answer these questions, we perform a comprehensive case study over 16 datasets spanning 9 languages, 9 domains, and 11 SOTA LLMs. We specifically focus on the following four research questions: (iHow well can human annotators distinguish human-written text from SOTA LLMs generations? (iiWhat are the notably detectable linguistic qualities in human- and AI-authored texts that shape decisions about text origin? (iiiCan a prompting strategy fill the gap between human and machine? and (ivDo humans always prefer human-authored text? Our comprehensive analysis exposes the major gaps between current LLMs and human expectations, suggesting possible directions for future LLM development. Our human detection evaluation challenges previous findings in terms of random-guess detection capability for humans. Moreover, our human preference collection can serve as a potential useful source for multilingual LLM alignment. Our contributions are as follows:
@@
  • Our extensive case study shows that expert annotators achieve detection accuracy of 87.6% on average over $\sim$9K examples, demonstrating that humans can identify LLM outputs and safeguard against LLM misuse. We identified a spectrum of language- and domain-specific distinguishable features that distinguish human-written from machine-generated text, and we summarized five generalized distinctions across languages, which clearly expose the limitations of current LLM outputs and indicate directions for improvements.
@@
  • Humans do not always like human-written text: preferences are evenly split between human vs. machine, with a strong tendency to favor machine-generated text (MGT) when annotators can not clearly identify the origin of a given text.
  • We release 17k original and 32k improved MGTs, labels of 13.5k human detection instances, 5k preferences and 1.6k fill-in-the-gap survey, and metadata for 19 annotators to facilitate future work on human vs. MGT detection, preference, and their relationship with individual characteristics.
@@
Shamardina et al. (2022, 2025) 2-class Russian 14 decoder-only, encoder-decoder models Wikipedia, news, social media, RNC, diaries, strategic documents 2,500 Toloka 66.66%
@@
Guo et al. (2023) 2-class English GPT-3.5 Wikipedia, QA 150 8 experts + 9 layman 48-90%
Guo et al. (2023) 2-class Chinese GPT-3.5 Wikipedia, QA 210 8 experts + 9 layman 54-93%
@@
Таблица 1: Experts refers to individuals who are frequent users of LLMs, laymen are people who have never heard of or used LLMs or have rarely used them, AMT is Amazon Mechanical Turk. RNC is the Russian national corpus.
@@ 2 Case Study
Previous studies presented the difficulty humans face in distinguishing SOTA LLM-generated content from human-written text, often resulting in a random guess (see more in Appendix A). However, most evaluations focused on English and generations by GPT-3.5-turbo, leaving the detectability of MGT in other languages and LLMs uncertain. To verify whether this observation can generalize to other languages and more advanced LLMs, we collected LLM generations based on 16 datasets across 9 domains and 9 languages. 19 native speakers who are either LLM researchers or practitioners performed human evaluations, investigating (i) whether expert humans can correctly discern human vs. AI outputs, and (ii) whether humans prefer fellow human answers or LLM responses.
@@ 2.1 Task and Dataset
The goal is to identify whether the text was written by a human or generated by models given a single text, or to recognize which text is written by a human given a pair of texts: one human-written and one machine-generated. In data collection, we focus on datasets from common domains such as community QA, news, tweets, and government reports, alongside domains requiring high-integrity LLM applications, including educational and academic contexts, such as accurate knowledge verification in Wikipedia-style texts, and identifying the authorship of student essays and peer reviews. For a given language and given a dataset, we sampled 300-600 human texts and then generated corresponding machine text using two SOTA LLMs: a multilingual model (e.g., GPT or Claude series) and a language-specific model (e.g., ChatGLM or Qwen for Chinese and AceGPT for Arabic), to analyze the impact of different LLMs on detection performance, particularly for non-English languages. Table 4 shows the statistical information of the 16 datasets, and Appendix C elaborates the generation prompts and collection details.
@@
IV. Triplet-Three-Class (hwt, mgt1, mgt2) Given a set of texts (text1, text2, text3), identify which one is human-written? One of the text1, text2 and text3 must be hwt, and others are mgt randomly sampled from $mgt_{i}$. A. text1; B. text2; C. text3 parallel human and multiple LLM generations are collected to make comparisons.
@@
2.2 Human Detection Setups
Annotation Settings
[…]
@@ 2.2 Human Detection Setups
We highlight that this study aims to explore the upper bound of human detection capability, rather than represent the full population distribution. Therefore, we did not recruit lay annotators unfamiliar with LLMs. Instead, we conducted in-house annotation with BSc., MSc., and PhD. students, as well as postdoctoral researchers, all with NLP expertise and prior experience with LLM-generated texts. Annotators were native speakers of the respective languages and completed annotations independently (see Appendix A.3 for details). While using expert annotators may limit the generalizability of results to lay users, our goal is to establish expert performance ceilings. If trained experts struggle to detect machine-generated text, it is reasonable to expect lower performance from the general public. Lay annotator studies are important for practical applications but do not substitute for upper-bound evaluation. Future work will include lay annotators to capture a broader distribution of human detection capabilities.
@@
Таблица 3: Human annotator detection accuracy over 16 datasets and 9 languages: we have a total of 30 annotation settings and 19 unique human annotators. The simple average accuracy of the human expert guesses is 87.6%.
3 Human Detection
We performed an extensive case study on 9k instances across 9 languages to verify how difficult it is for native speakers to detect AI outputs in everyday domains. Table 3 demonstrates that human detection accuracy can reach 87.6% on average. This reveals that this is not particularly difficult for native human experts, contrary to what previous studies have reported. Below, we zoom into the impact of various factors.
@@ 3.3 Generator
The impact of generators is associated with languages and domains. While there are minimal differences for Chinese, where accuracy is 100% vs. 98% for GPT-4o vs. Qwen-turbo, there are sizable differences for Italian and Arabic. Based on Italian DICE News, the same annotator detected generations by Anita (an Italian fine-tuned Llama3-8B), Llama3-405B, and GPT-4o, achieving accuracy of 88%, 99%, and 100%, respectively. Similarly, for Arabic tweets, GPT-4o’s outputs are more similar to human text and thus more difficult to detect compared to those by Qwen2, as shown in Table 6.
@@ 3.4 Annotation Setting
We conducted the majority of our annotations under setting I. pair-binary: given a pair (hwt, mgt), it asks to identify which of the two texts is human-written. We assumed that more complex settings would be more challenging. For instance, II. pair-four-class should be harder, as each of the texts could be human- or machine-generated, independently of the other. Yet, for Chinese student essays, the performance for II does not degrade compared to I. Similarly, for government reports in IV. triplet-three-class, where the annotators have to select the human-written text among three candidates, there was no degradation compared to I. However, III. single-binary proves to be more difficult than I. pair-binary for both Arabic and Russian. While domain differences do exist, e.g., tweets vs. summary vs. news in Arabic, and news vs. summary for Russian, the substantial performance gap (>20%) can still be partially attributed to the annotation settings. Overall, comparing the results for I. vs. III., it is easier to distinguish machine-generated content when given a comparison pair, rather than for single answer. Yet, introducing text triplets or increasing the number of machine-generated or human-written texts had minimal impact on detection performance. Moreover, using few shots before detection boosted the confidence of the annotators, resulting in higher accuracy compared to zero-shot. For datasets with high accuracy, before seeing labeled samples, the annotators found the distinction to be obvious. After seeing a few examples, the annotator was extremely confident in distinguishing human vs. machine text based on indicative features of MGT.
@@ 3.6 Distinguishable Factors
We summarize five remarkable distinguishable signals between machine-generated vs. human-written text across the 16 datasets and the 9 languages. Appendix D provides detailed distinctions between hwt and mgt for each language and dataset, where valuable expert insights into the limitations and gaps of current LLM outputs compared to human-written text can inform future model improvements.
@@
  • Machine-generated text shows a mixture of other languages. Non-English language responses often contain some English parts, and sometimes mix other languages, e.g., Japanese responses contain Chinese or Korean, particularly for less capable models. This is very rare in human text.
@@ 4 Can Prompting Fill in the Gap?
Given that LLMs can closely follow instructions and their outputs are heavily influenced by the system and the user prompts, we investigated whether explicitly instructing them to mimic human style can help narrow the gap. Responding to the distinguishable factors summarized in Appendix D for each dataset, we asked the human annotators to craft new prompts, aiming to improve the generations and to reduce the gap between human-written and LLM-generated texts. This involved trying instructions that (i) incorporate specific details and references, (ii) avoid formulaic structures and formats, e.g., bullet points and Markdown, and (iii) generate texts of varying length, structure, and sentiment. Table 8 presents both the original and the improved prompts for all datasets. Measurements: We re-generated the machine-generated parts of the text pairs, using the same models with improved prompts, and then sampled 200–600 examples from each dataset to assess whether and to what extent, the prompting strategy narrowed the gap between human-written and machine-generated texts. We used two approaches: (i) intuition-oriented fill-the-gap survey, where the original annotators evaluated whether the newly-generated text bridged the gap (Yes, No, or Partially), and (iidownstream detection, where we compared the detection accuracy before and after applying the improved prompts. A decline in detection accuracy indicated a reduced distinction between human and machine text, making the differentiation more difficult, and further revealing that prompting was effective. While the survey is inherently slightly subjective, it remains valuable, as human intuition plays a crucial role in evaluating the perceived quality of model outputs. By combining this intuition-driven assessment with quantitative detection results, we aim to draw more robust and convincing conclusions.
@@
The original annotators who conducted the detection on earlier generations were asked to evaluate whether the new prompts addressed the gaps for each example: Yes, No, and Partially. The distributions across the six datasets in Figure 1 shows that in about 50% of the cases, the prompt adjustments were effective to either fully or partially mitigate the gaps. Large improvements were observed for Kazakh Wikipedia and Arabic tweets. For the former, the revised prompt reduced repetitive sentence patterns (more diverse), but the formulaic expressions were not entirely eliminated. New outputs also included more concrete information, such as dates and names, while the inclusion of culturally-nuanced details remained challenging. The newly-generated Arabic tweets could touch on relatable human topics and express genuine emotions tied to daily experiences; however, the frequently added irrelevant hashtags at the end of the tweets made them easily identifiable as machine-generated. Moreover, the tweets often leaned on an overly optimistic tone even when negative experiences were mentioned. The annotators for English peer reviews noted that, despite the prompt adjustments, the models remained highly formulaic in their outputs, the length of the reviews remained relatively uniform, and the overall structure lacked variance. This may be due to the inherent nature of the peer review domain, while human reviews exhibited more variability in both length and structure, feeling more organic. Similar issues remained for Chinese student essays (formulaic structure by using ‘‘firstly, then, moreover, finally’’ persisted) and government reports (certain repetitive phrases). See more details in D.2. Overall, adjusting the prompts did fill some gaps, but cultural nuances, diversity of length, structure and phrases, and sentiment adaption to scenarios remained challenging. See more in Appendix E. With the intuitive survey results, we further conducted a second round of MGT detection on the new generations for objective results, involving both human evaluation and automated detection.
Рис. 2: Human detection accuracy dropped from 87.6 to 72.5 for the original vs. the improved generations.